A small block with a mass of 0.0400 kg is moving in the π‘₯𝑦-plane. The net force on the block is described by the potential energy function π‘ˆ(π‘₯) = (5.80 𝐽 π‘š2 ⁄ )π‘₯ 2 βˆ’ (3.60 𝐽 π‘š3 ⁄ )𝑦 3 . What are the magnitude and direction of the acceleration of the block when it is at the point (π‘₯ = 0.300π‘š, 𝑦 = 0.600π‘š)?Β 


1
Expert's answer
2021-10-31T18:11:08-0400

Find the energy at the given points:


"U_x(0.3)=5.8·0.3^2=0.522\text{ J},\\
U_y(0.6)=3.6·0.6^3=0.778\text{ J}."

The force can be fund as


"F=U\/r."

The acceleration is


"a=F/m=U(rm),\\\space\\
a_x=U_x/(xm)=0.522/(0.3·0.04)=43.5\text{ m/s}^2,\\
a_y=0.778/(0.6·0.04))=32.4\text{ m/s}^2.\\
a=\sqrt{a_x^2+a_y^2}=54.2\text{ m/s}^2,\\\space\\
\theta=\arctan\frac{a_y}{a_x}=36.7°"

from positive x-direction to positive y.